OpenStudy (kimberly_pr):

Will Medal! Given an arithmetic sequence in the table below, create the explicit formula and list any restrictions to the domain. n an 1 40 2 47 3 54

OpenStudy (kimberly_pr):

Answer choices: an = 40 + 7(n − 1) where n ≥ 40 an = 40 + 7(n − 1) where n ≥ 1 an = 40 − 7(n − 1) where n ≥ 40 an = 40 − 7(n − 1) where n ≥ 1

OpenStudy (kimberly_pr):

@ganeshie8 @jigglypuff314

OpenStudy (mathmale):

What's the "common difference?" In other words, what number separates 47 from 40, and which number separates 54 from 47?

OpenStudy (dayakar):

what is the first term a = what is the common difference = a2-a1 = can u find these?

OpenStudy (kimberly_pr):

7 is the common difference

OpenStudy (kimberly_pr):

And the first term is 40, right?

OpenStudy (kimberly_pr):

S what do I do now?

OpenStudy (kimberly_pr):

*so

OpenStudy (dayakar):

very good a= 40 , d= 7 an = a + (n-1) d plug in above values

OpenStudy (kimberly_pr):

ok so the equation would be 40n= 40 + (n-1)7 ?

OpenStudy (dayakar):

no ,here an means nth term an = 40 + (n-1)7 = 40 +7(n-1) and \[n \ge 1\]

OpenStudy (dayakar):

n is number of terms are always 0 or poisitive

OpenStudy (kimberly_pr):

Ok, so then the answer is an= 40 + (n-1)7 where n is greater than or equal to 1 Thank you sooo much for all your help and your clarification :)